OpenStudy (anonymous):

Let S be subspace of R^3 spanned by x=(1,-1,1)^T a) Find a basic for S^ʇ

OpenStudy (anonymous):



OpenStudy (anonymous):

you want to find the null space for the 1 x 3 matrix [ 1 -1 1].

OpenStudy (anonymous):

or rather, you want to find a basis for the null space.

OpenStudy (anonymous):

so we find the null space then find the basic of that null space

OpenStudy (anonymous):

yeah. Its a little odd since the matrix is 1 x 3, but it will work out the same.

OpenStudy (anonymous):

Think of it this way, we know two vectors are perpendicular when their dot product is zero. So the vector (x,y,z) is perpendicular to (1,-1,1) when x-y+z=0. But this is the same as solving the matrix equation:\[A\vec{x}=0\]when A is the matrix [ 1 -1 1]

OpenStudy (anonymous):

so x be any vector that makes right side equal to 0

OpenStudy (anonymous):

like [0 0 0] ?

OpenStudy (anonymous):

while that is true, you are looking for non-zero vectors. The zero vector will always be a solution to Ax = 0, thats called the trivial solution.

OpenStudy (anonymous):

since we have 1 variable but 3 equation; wouldn't it likely be inconsistent?

OpenStudy (anonymous):

Since there is only one equation, but three variables, we can make two of the variables free, and see what the last one has to be.

OpenStudy (anonymous):

1 x1 - x2 +x3=0 x1= a x2=b a-b+x3=0 -a+b=x3

OpenStudy (anonymous):

thats the right idea. Just make them numbers now so we can grab a basis. Like x1=1 and x2=0, or something like that.

OpenStudy (anonymous):

There is a more straightforward way of doing this, which i'll post in a bit, but this isnt bad either.

OpenStudy (anonymous):

x1=1 x2==0 -a+b=x3 -1+0=x3 x3=-1

OpenStudy (anonymous):

So the vector (1, 0, -1) is one of your basis vectors. Now to get another that is independent of this one, make x1 = 0, and x2 = 1. See what x3 is, and that will be the second.

OpenStudy (anonymous):

x1=0 x2=1 1=x3 0 1 1

OpenStudy (anonymous):

that works. As you can see, the vector (1,-1,1) is perpendicular to both of the vectors you obtained, so it will be perpendicular to any linear combination of those vectors, ie the span of them. That subspace will be perpendicular to the span of (1,-1,1).

OpenStudy (anonymous):

Here is another method. Even though the numerical value of the vectors are different, they do span the same space. So no worries.

OpenStudy (anonymous):

quick question what does that upside down T next to S mean?

OpenStudy (anonymous):

O.o i was assuming it meant perpendicular <,<

OpenStudy (anonymous):

or orthogonal, like this: http://en.wikipedia.org/wiki/Orthogonal_complement

OpenStudy (anonymous):

yes , orthogonal complement

OpenStudy (anonymous):

oh ok, we are still good then lol.

OpenStudy (anonymous):

thank you so much, joe; helpful like always

OpenStudy (anonymous):

Linear Algebra rules >.> lolol. yw!